OpenStudy (anonymous):

Find the area of the triangle with base 2_4^3 yards and height 4^5 yards. Area of the triangle is equal to half of the product of it’s base and height.

OpenStudy (anonymous):

\[2\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

4^9

OpenStudy (anonymous):

Find the area of the triangle with base 2_3^4 yards and height 4^9 yards. Area of the triangle is equal to half of the product of it’s base and height.

OpenStudy (anonymous):

Is the height really 4 raised to the 9th power?!? That's what 4^9 means. Or did you mean the fraction 4/9

OpenStudy (anonymous):

4/9

OpenStudy (anonymous):

2 3/4

OpenStudy (anonymous):

Ok\[2\frac{ 3 }{ 4 }=\frac{ 11 }{ 4 }\] Do you get that? You want to convert everything to normal fractions first so now it's just\[\frac{ 1 }{ 2 }\times \frac{ 11 }{ 4 }\times \frac{ 4 }{ 9 }\]

OpenStudy (anonymous):

The 4's cancel there so it's just\[\frac{ 11 }{ 2\times9 }=\frac{ 11 }{ 18 }\]

OpenStudy (anonymous):

ok i came up with 12/15

OpenStudy (anonymous):

\[\frac{ 11 }{ 18 }yds ^{2}\]

OpenStudy (anonymous):

7/18 was answer

OpenStudy (anonymous):

Are you sure the base isn't one and 3/4ths rather than 2 and 3/4ths. If you use 1 and 3/4ths as the base you get 7/18. That's all I can think of.